OpenStudy (anonymous):

Simple question: Is this a direct variation? Age x.... 10, 11, 12, 13 Grade y.... 5, 6, 7, 8 (that's supposed to be a table) I do not think it is... that's correct, right? You don't get the same thing when you divide 10 and 5 and when you divide 13 and 8

OpenStudy (opcode):

Since both the x and y values increase, then this is direct variation. Direct variation is written as: \[y = kx\] Where k is the constant of variation. \[5 = k(10)\] \[k = \frac{1}{2}\] \[y = \dfrac{1}{2}x\] What's the issue?

OpenStudy (campbell_st):

nope... for direct variation... you need y = kx each value of x is multiplied by the same value...

OpenStudy (campbell_st):

if it was direct variation the grade values would be 1/2 x 10 = 5 1/2 x 11 = 5.5 1/2 x 12 =6 etc